OpenStudy (anonymous):

Sam stores manure for his farm in a rectangular box. The length of the box is 7 feet, width is 9 inches, and height is 6 inches. Each bag of manure he bought contains 2.625 cubic feet of manure. How many bags of manure will fill the box completely? I think the answer is: 1 bag of manure will fill the box completely.

OpenStudy (perl):

this problem has different units, feet and inches, so be careful , and good luck :)

OpenStudy (perl):

9 inches = 9/12 feet = 3/4 foot 6 inches = 6/12 feet = 1/2 foot so we are given the volume of the box is V = 7*3/4 * 1/2 = 2.625 cubit feet

OpenStudy (perl):

so the answer is 1 bag of poo

OpenStudy (anonymous):

Thank you very much! :) I'm surprised I got it right hehe.

OpenStudy (perl):

medal please :)